OpenStudy (anonymous):

a rectangle has an area of 120cm^2. its length is 2cm longer than its width. Use a parabola and quadratic to find the dimensions??

OpenStudy (radar):

\[(w)(w+2)=w ^{2}+2w=120\]\[w ^{2}+2w-120=0\] (w+12)(w-10)=0 w=10 length=w+2=10+2=12

OpenStudy (radar):

The quadratic equation is a parabola

OpenStudy (anonymous):

thanks